Circular No. 5910

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

COMET McNAUGHT-RUSSELL (1993v) Robert H. McNaught reports the following measurements of a comet

that he discovered on a U.K. Schmidt Telescope plate taken by Kenneth S. Russell. The comet shows a coma of diameter 8′′ and a very faint tail 10′′ long in p.a. 20^{o}. The object’s westward motion was confirmed in an observation with the Siding Spring 0.61-m reflector (+ CCD) on Dec. 18.

1993 UT  α_{2000}  δ_{2000}  m_{1}
Dec. 17.46003  3^{h}06^{m}33.^{s}47  −56^{o}33′24′′.0  17.5
17.50865  3 06 28.62  −56 33 23.1

NOVA CASSIOPEIAE 1993 Visual magnitude estimates indicate a brightening of this object: Dec.

15.13 UT, 6.4 (A. Hale, near Orogrande, NM); 16.12, 6.4 (C. Hergenrother, Tucson, AZ); 17.11, 6.3 (C. E. Spratt, Victoria, BC); 17.23, 5.6 (Hale, El Paso, TX); 17.712, 5.8 (P. Rapavý, Rimavská Sobota, Slovakia); 18.08, 5.3 (C. E. Spratt, Victoria, BC); 18.11, 5.6 (Hergenrother).

S. R. Trammell and R. A. Benjamin, University of Texas, report: “Optical spectropolarimetry (range 450-720 nm; dispersion 0.269 nm/pixel) obtained with the 2.7-m telescope at McDonald Observatory on Dec. 14.06 UT, shows a high-continuum polarization of 1.84 ± 0.05 percent at position angle 61^{o} ± 1^{o}. There is a 0.4-percent drop in polarization across Hα with a position angle rotation of 10^{o} ± 2^{o}, which indicates that some polarization is intrinsic to the nova and suggests an asymmetric geometry.”

C. E. Woodward, University of Wyoming; and M. A. Greenhouse, National Air and Space Museum, write: “Velocity-resolved spectroscopy of N Cas 1993, obtained on Dec. 15.2 UT using the NASA Infrared Telescope Facility (+ CSHELL), revealed emission from hydrogen Brγ with FWZI ∼ 1680 km/s. Substantial structure was observed in the emission-line profile. Infrared photometry obtained by G. F. Lawrence and R. D. Gehrz, University of Minnesota, using the Mt. Lemmon 1.5-m telescope resulted in the following broadband magnitudes: Dec. 14.1 UT, J = 4.82 ± 0.03, H = 4.48 ± 0.03, K = 4.18 ± 0.03, L′ = 3.65 ± 0.03, M = 3.00 ± 0.03 (with InSb photometer); Dec. 15.1, H = 4.55 ± 0.03, K = 4.12 ± 0.03, L = 3.92 ± 0.03 (with bolometer). The observed color magnitudes are consistent with the nova being in a free-free evolutionary state at this epoch.”

1993 December 18 Daniel W. E. Green